Jen is stockpiling his treats in a hole. He sets a 500g box of treats on a ramp angled at 65 degrees. The core didn’t of friction is 0.15. 

  1. how fast will the box accelerate down the ramp 
  2. If the ramp is 1.5m, how long will it take the box to get to the bottom 
1
Expert's answer
2021-07-15T10:51:38-0400
"a=g(\sin{65}-\mu\cos{65})\\a=9.8(\sin{65}-0.15\cos{65})=8.3\frac{m}{s^2}\\"

"l=0.5at^2\\1.5=0.5(8.3)t^2\\t=0.6\ s"